OpenStudy (anonymous):

a triangle's base is 1/3 the height and the area is 6. find the base of the triangle

OpenStudy (ddcamp):

If base is one third the height, we can also say that the height is three times the base, or h=3B Plugging this into the formula for the area of the triangle, we have: A=½B×h A=½(B)(3B) A=(3/2)b² We're told that the area is 6, so now you just need to solve: 6 = (3/2)B²